a) Consider the following functional relation, f, defined as:


f : R \rightarrow R, f(x) = x2

Determine whether or not f is a bijection. If it is, prove it. If it is not, show why it is not.




b) Consider the set


F = {y | y = ax3 + b},


a, b being constants such that a \ne 0 and x \in R.
Is F equivalent to R? If so, prove it. If not, explain in details why it is not the case.
1
Expert's answer
2022-01-31T14:08:49-0500

a.)

"f:\R \to \R, f(x)=x^2"

This function is not a bijection since it is not one-to-one. We can see this as follows:

"f(-2)=f(2)=4", but "-2 \neq 2". Thus "f" is not injective. Hence it is not a bijection.


b.)

"F=\{y|y=ax^3+b\}"

Yes, "F" is equivalent to "\R" since will can always find a bijection between the two set.

"f:F\to \R"